OpenStudy (anonymous):

In degree measure the ngles pie/3 pie/4 and pie/2 are respectively

OpenStudy (anonymous):

multiply each by 180/pie 60,45,90 degrees

Parth (parthkohli):

\( \color{Black}{\Rightarrow 2\pi = 360^{\circ} }\) \( \color{Black}{\Rightarrow \pi = 180^{\circ} }\) So, \( \color{Black}{\Rightarrow \pi \div 2 = 90^{\circ} }\) \( \color{Black}{\Rightarrow \pi \div 3 = 60^{\circ} }\) \( \color{Black}{\Rightarrow \pi \div 4 = 45^{\circ}}\)